The 1000-turn, 0.4 m diameter coil of a generator is rotated at 120 rpm in a 0.038 T field. In order
to be able to measure the emf induced by the generator up to the maximum, a galvanometer of
10
1
Expert's answer
2020-08-02T15:09:14-0400

As per the given question,

Number of turn"(N)=1000"

Diameter of the coil "(d)=0.4m"

Frequency of the rotation of the coil "(\omega)=2\pi \times 120\times \frac{1}{60} =4\pi Hz"

Magnetic field"(B)=0.038T"

"E=NBA\omega"

Now, substituting the values in equation,

"E=1000\times 0.038\times \pi (\frac{0.4}{2})^2\times 4\pi Volt"

"=1000\times 0.038\times \pi\times 0.04 \times 4\pi Volt"

"=19.1 Volt"